The sum of the ages of Mr.Lin and his son is 38 years. In three years' time, Mr.Lin will be three times as old as his son. Find thier present ages.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)
Answer by ikleyn(52804)

.
The sum of the ages of Mr.Lin and his son is 38 years. In three years' time, Mr.Lin will be three times as old as his son. Find thier present ages.
~~~~~~~~~~~~~~~~~~~~~~~~

Let L = Mr.Lin's present age,
    S = his son's present age.

First equation  is

L + S = 38.       (1)

The second equation is 

L + 3 = 3*(S+3)   (2)   ("In three years' time, Mr.Lin will be three times as old as his son.")

From (1), express S = 38-L and substitute it into (2). You  will get

L + 3 = 3*((38-L)+3),   or

L + 3 = 114 - 3L + 9,

4L = 114 + 9 - 3, 

4L = 120,

L = 30.

Answer.  Mr.Lin's present age is 30 years now. His son is 8 years old now.

For many other solved age problems see the lessons
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.